The lengths of the sides of a rectangle are in the ratio 5 : 3. The perimeter is 32 cm. What are the dimensions?
Answer by Alan3354(69443)

L + W = 1/2 the perimeter = 16 cm
5+3 = 8
16/8 = 2
-------
L = 5*2 = 10 cm
W = 3*2 = 6 cm